in 1919, the winner of a competition was paid $100. in 2019, the winner’s prize was $235,000. what will the winner’s prize be in 2040 if the prize continues increasing at the same rate ?
Answer #1

1. Information provided:

Present value= $100

Future value= $235,000

Time= 2019 – 1919= 100 years

The interest rate is computed by entering the below in a financial calculator:

PV= -100

FV= 235,000

N= 100

Press the CPT key and I/Y to compute the interest rate.

The value obtained is 8.0714.

Therefore, the interest rate is 8.0714%.

2. Information provided:

Present value= $235,000

Time= 2040 – 2019= 21 years

Interest rate= 8.0714%

The question is solved by computing the future value.

Enter the below in a financial calculator to compute the future value:

PV= -235,000

N= 21

I/Y= 8.0714

Press the CPT key and FV to compute the future value.

The value obtained is 1,199,483.257.

Therefore, the winner’s prize in 2040 will be $1,199,483.26.

In case of any query, kindly comment on the solution.